Find the quotient and remainder of f(x)=x^4-2 divided by p(x)=x-1.
Answer by stanbon(75887)

Find the quotient and remainder of f(x)=x^4-2 divided by p(x)=x-1.
Using synthetic division I get the following:
1 ) 1 0 0 0 -2
1 1 1 1 | -1
This shows the quotient is x^3+x^2+x+1
with remainder -1
Cheers,
Stan H.